OpenStudy (anonymous):

1-2sin^2 2theta rewrite with double angle identity

OpenStudy (raden):

\[1 - 2\sin^{2} x = \cos(2x)\]

OpenStudy (raden):

sorry, that should be theta, not x :)

OpenStudy (anonymous):

[1-2\sin ^{2}2\theta\] Hi thanks for the help! yeah I know that's the trig identity that I should use but this is the equation so it's not just a basic trig identity

OpenStudy (anonymous):

\[1-2\sin ^{2}2\theta\]

OpenStudy (raden):

\[1 - 2\sin^{2} 2\theta \] = \[1 - \sin^{2} 2\theta - \sin^{2} 2\theta \] = \[\cos^{2} 2\theta - \sin^{2} 2\theta \] = \[\cos(2(2\theta)) \] = \[\cos(4\theta)\]

OpenStudy (anonymous):

so you treated the identity as though there wasn't a 2 before the theta? sorry just trying to comprehend...THANK YOU SO MUCH

OpenStudy (raden):

just see wrong at the first, i am sorry welcome